OpenStudy (anonymous):

suppose sin u= x and u is between 0 and pi/2. what is sec u?

OpenStudy (zarkon):

\[\sec(u)=\frac{1}{\cos(u)}\] and \[\cos(u)=\sqrt{1-\sin^2(u)}\]